OpenStudy (anonymous):

Find the derivative of y=arcsin(x^2)+ arcsin(sqrt(1-x^4))

OpenStudy (anonymous):

\[y=\arcsin x^{2} + \arcsin (\sqrt{1-x^{4}})\]

OpenStudy (anonymous):

\[\frac{ d }{ dx } arcsinu = \frac{ 1 }{ \sqrt{1-u^{2}} } \frac{ du }{ dx }\]

OpenStudy (anonymous):

Am I on the right track? \[\frac{ 1 }{ \sqrt{1-x^{4}} } \frac{ d(x^{2} }{ dx }+ \frac{ 1 }{ \sqrt{1-\sqrt{1-x^{4}}} } \frac{ d(\sqrt{1-x^{4}}) }{ dx }\]